In a fort, ration for 2000 people was sufficient for
54 days. After 15 days, more people came and the
ration lasted only for 20 more days. How many people
came?

Answer by richwmiller(17219)

54*2000 food days
54*2000-2000*15=39*2000
39*2000 food days
39*2000/20=3900 people
3900-2000=1900 more people arrived